OpenStudy (kg1975):

Plz Help @Taylor0402

OpenStudy (kg1975):

@Taylor0402

OpenStudy (taylor0402):

ok so were were we?

OpenStudy (kg1975):

uhm let me find out

OpenStudy (kg1975):

last thing u said math wise is.... if any number has a variable next to it (a letter) then you ALWAYS multiply unless it says to do otherwise

OpenStudy (taylor0402):

yes so if it says in your equation 2x-y then you would do 2times3+3

OpenStudy (taylor0402):

that minus sign is supposed to be a plus

OpenStudy (kg1975):

2*3+3? ohhhhhhhhhhhhhhhh now i get it! a lil bit

OpenStudy (taylor0402):

yayyyyy

OpenStudy (taylor0402):

so what would the bottom equation be

OpenStudy (kg1975):

What is the value of the expression \[\frac{ 4x-y }{ 2y+x }\] when x = 3 and y = 3?

OpenStudy (kg1975):

okie i remember u plugged the numbers in... \[\frac{ 4*3-3 }{ 2*3+3 }\]

OpenStudy (taylor0402):

good thats correct now do you know what to do?

OpenStudy (kg1975):

P : Parenthesis E : Exponents M: Multiply D: Divide A: Addition S : Subtraction So we multiply first right?

OpenStudy (kg1975):

4*3?

OpenStudy (taylor0402):

ye and

OpenStudy (kg1975):

2*3?

OpenStudy (taylor0402):

very good so what does the equation look like now?

OpenStudy (kg1975):

4*3=12 and 2*3=6 \[\frac{ 12-3 }{ 6+3 }\]

OpenStudy (kg1975):

right?

OpenStudy (taylor0402):

perfect now we have 2 more steps so then we solve the 2 equations

OpenStudy (kg1975):

ill be right back while u type i have to use the restroom xD

OpenStudy (taylor0402):

so 12-3+9 and 6+3=9

OpenStudy (taylor0402):

and because of that \[9 \over 9\] the bar in the middle means to divide the two equations and 9 divided by 9 is 1

OpenStudy (kg1975):

im back

OpenStudy (kg1975):

wait why do u divide the fraction?

OpenStudy (unofficialllyy):

@Taylor0402 Please do not give out answers.

OpenStudy (kg1975):

:T

OpenStudy (taylor0402):

so the finial answer is 1

OpenStudy (taylor0402):

i didnt give it i explained it

OpenStudy (kg1975):

PLZ STOP! SHEESH!!!

OpenStudy (kg1975):

Ty Taylor that makes alott more sense!!

OpenStudy (taylor0402):

no problem hey do you have a snapchat?

OpenStudy (kg1975):

i wish... but no i dont have any of those things like kik or snapchat or skype DX

OpenStudy (kg1975):

My parents wont let me...

OpenStudy (kg1975):

idk why

OpenStudy (kg1975):

BTW i gave u the medal :D

OpenStudy (taylor0402):

man that kind of stinks well i have gtg help more ppl so have a good day :) thanks

OpenStudy (kg1975):

ty!